# Position: A taxonomy for reporting and describing AI security incidents

### v1

@@ Appendix A Definitions
AI model. An AI model is a machine-based system that, for explicit or implicit objectives, infers from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions, as part 1 of the OECD definition [^14] https://oecd.ai/en/wonk/ai-system-definition-update.
@@
Meta Information 1.1 Incident ID Incident ID The unique identifier of the incident. ID_1716
@@
1.2 Incident Reporter Name of incident reporter Name or also username from within recording. John Doe / AIIncidentreporter001
1.3 Reporting Organization Name of reporting organization The legal name of the reporting organization. Acme Inc.
1.4 Incident Date Date the incident ocurred Date, Time interval, if unkown estimate interval 13/12/2024
@@
AI System Information 2.1 Model type Specify the ML model type. Generative or predictive AI? prediction, generative, but also more concrete, Transformer, DNN, ConvNet, SVM etc Support Vector Machine
@@
2.2 Model version High level identifier, as noted in model card. This identifier should allow for correct identification of the specific model version. Llama3.2:1B
@@
2.3 Model specification Provide as precise information on the model as possible. Information could contain URLs to or files with weights, full model files or model cards. ’URL’
@@
2.4 Lifecycle stage Define the stage of the AI lifecycle which the model was in when the incident occurred. Stages of the AI lifecycle are: Inception, Design & Development, Verification & Validation, Deployment, Operation and Monitoring, Re-evaluate, Retirement Deployment
@@
2.5 Domain Did the incident occurr in a physical environment? Did the incident affect an AI that is interacting with physical objects or otherwise physically with humans? Yes, pedestrian not detected.
@@
2.6 Intended use Described the intended use of the AI system. Intended purpose means the use for which an AI system is intended by the provider, including the specific context and conditions of use, as specified in the information supplied by the provider in the instructions for use, promotional or sales materials and statements, as well as in the technical documentation. Object detector to tag cats on image, Chatbot to reply to questions about air plane fares, etc
@@
2.7 Actual use during incident, if differing from 2.6 Describe the use of the AI system during the incident, if differeing from intended use (2.6) This may mean the usage for which an AI system was not intended by the provider, including the specific context and conditions of the actual use. Used to tag foxes instead of cats, used to get legal advice model was not trained for.
@@
2.8 Model supply chain - Data Was publicly available data used? If yes, from where, was data extended or adapted and if yes how? Specify dataset with name and URL, as well as any modifications that the data underwent before being used. For example: Model used MNIST data from ’URL’, where we added 1000 images that were handlabelled by our team.
@@
2.9 Model Supply Chain - Model weights Are the model weights derived from an existing model, and if yes, which one and how were they adapted? If the original design or weights have a name, provide this name or possibly a URL. Further describe all steps the model undewerwent to adapt it to the current status. For example: Based on Llama3.2:1B, but finetuned on the above described dataset for n iterations using optimzier o and loss l.
@@
2.10 Model Supply Chain - Code What public code was used to generate and run the model? Provide libraries used with their version name, and also ennumnerate github repositories with link and access data or other re-used code from the project. For example: Pytorch version 2.4.1. and Code from github repo ’this url’ was used, forked January 1st 2024.
@@
2.11 System prompt If the system is a generative AI system, provide the system prompt. If known, report the entire system, prompt as used in the setting where the incident occurred. For example: ’please be polite to the user and don’t hallucinate. The user writes: ’
@@
Table 3: Reporting scheme with description for annotators and examples, second third (Part 3.1 - Part 3.9).
@@
Incident specification 3.1 Materialization status Has the incident been reported prior or after its materialization? Was the attack found during red-teaming or scientific analysis, or as part of an exploited breach? Red teaming
3.2 Root cause Describe the root cause of the incident. What is the underlying reason, design choice, or property of the system that enables the attack? For example: ’Attacker relied on infinity norm attack the model wasn’t tested for’, or ’API allowed infinite amount of queries, hence enabling crating a copycat training dataset’
3.3 Malicious Intent Is there evidence that an entity or a person purposely caused the incident by targeting the AI systems security? Is the incident an example of an attempt to destroy, expose, alter, disable, steal or gain unauthorized access to or make unauthorized use of the AI system? For example: ’Yes, the attacker altered the sample until it passed the AI security check’ or ’No, the user did not know that the system was not trained on this specific subset of the training data (Disease B). Hence, the system provided a wrong output.’
3.4 Attack type Specify the concrete type of attack that has been carried out. Any attack framework, tagging based on the attacks in Sect. 2.3 Poisoning
3.5 Target asset Did the attack target the model explicitly and the model only? Based on the definition in Sect. 2.1 Yes, model was targeted.
3.6 Hardware What hardware was used, and could it have affected the incident? Please describe the specific hardware used on the device or server the application ran, and whether the setting implies that the hardware played a role in the incident. Mac M1, could not have affected the incident OR could have affected the incident via a glitch.
3.7 Access Which access was given to whom for model, inputs, training, test or evaluation data? Specify which levels of access were given to which users regarding each component. Training and evaluation data as well as model was only available to internal model team. The model could be queried via API by the client, but access control prevented any other access by 3rd parties to the model.
3.8 Security measures What security measures where used? Where they circumvented and if yes how? Give a list of specific security measures that were used including their configurations. Adversarial training using FGSM with epsilon 0.1, 0.2 and 0.3 was used on randomly selected 10% of the dataset were used. Defense was not circumvented as attack was based on L2 norm, not $L_{\infty}$ norm.
3.9 Fail safe measures Where there any fail-safe measures implemented? Describe any fail-safe measures, such as rebooting the system, not outputting a classification, etc. System fell back on human intervention to prevent further harm.
Table 4: Reporting scheme with description for annotators and examples, last third (Part 4.1 - Part 5.1).
@@
Incident Implications 4.1 Regulatory damage List specific regulatory requirements that are violated due to the incident. Please specify any concrete legislation, standard or policy that has been violated by the incidents’ occurrence. European AI Act
4.2 Resulting harm What harm did the incident cause to any participating stakeholder? Did the incident lead to any of the following types of harm? Affecting physical health, inducing financial loss, damaging a physical object, affecting intangible property, affecting infrastructure, affecting natural environment, violation of human rights, civil liberties, civil rights or democratic norms. Harm to physical property of all users, mainly food but also some devices due to malfunction of the fridge AI.
4.3 Security Impact Describe how the incident affected the confidentiality, integrity, availability of the AI system. Describe whether the incident caused any leakage of confidential data, lead the system to output a wrong result or to exhibit a bias (integrity), or lead to the system being down or otherwise unusable (availability). During the incident, samples were manipulated to be classified as vehicles and thus harmed integrity.
4.4 Generalization Does the incident potentially affect other AI systems? Incident descriptions should allow to determine for a third model if a vulnerability in that model is likely. Other AI systems might also be affected if relying on the same dataset ”xyz”.
Appendix 5.1 Links to other incidents Links to similar / same cases in other /same database(s) In case you are aware of a incidents that has been reported in another database or that is similar to this incident in this database For example, this incident has had similar consequences like AIID incidents 38 or simply ’URL’
5.2 Comments Additional comments. Anything you’d like to add which has not been covered in the previous fields. For example, the attack does not work on model ABC1.2 and ABC1.4, but only on ABC1.3

### v2

@@ Appendix A Definitions
To distinguish AI system and AI model, we rely on the reasoning of the OECD (a): “Although different interpretations of the word “model” exist, in this document, an AI model is a core component of an AI system used to make inferences from inputs to produce outputs.”, referring to the following technical definition:
@@
AI model. Mathematical construct that generates an inference or prediction based on input data or information, from ISO/IEC 22989:2022.
@@
Meta Information T1.1 Incident ID Incident ID The unique identifier of the incident. ID_1716
@@
T1.2 Incident Reporter Name of incident reporter Name or also username from within recording. John Doe / AIIncidentreporter001
T1.3 Reporting Organization Name of reporting organization The legal name of the reporting organization. Acme Inc.
T1.4 Incident Date Date the incident ocurred Date, Time interval, if unkown estimate interval 13/12/2024
@@
AI System Information T2.1 Model type Specify the ML model type. Generative or predictive AI? prediction, generative, but also more concrete, Transformer, DNN, ConvNet, SVM etc Support Vector Machine
@@
T2.2 Model version High level identifier, as noted in model card. This identifier should allow for correct identification of the specific model version. Llama3.2:1B
@@
T2.3 Model specification Provide as precise information on the model as possible. Information could contain URLs to or files with weights, full model files or model cards. ’URL’
@@
T2.4 Lifecycle stage Define the stage of the AI lifecycle which the model was in when the incident occurred. Stages of the AI lifecycle are: Inception, Design & Development, Verification & Validation, Deployment, Operation and Monitoring, Re-evaluate, Retirement. Deployment
@@
T2.5 Domain Did the incident occurr in a physical environment? Did the incident affect an AI that is interacting with physical objects or otherwise physically with humans? Yes, pedestrian not detected.
@@
T2.6 Intended use Described the intended use of the AI system. Intended purpose means the use for which an AI system is intended by the provider, including the specific context and conditions of use, as specified in the information supplied by the provider in the instructions for use, promotional or sales materials and statements, as well as in the technical documentation. Object detector to tag cats on image, Chatbot to reply to questions about air plane fares, etc
@@
T2.7 Actual use during incident, if differing from 2.6 Describe the use of the AI system during the incident, if differeing from intended use (2.6) This may mean the usage for which an AI system was not intended by the provider, including the specific context and conditions of the actual use. Used to tag foxes instead of cats, used to get legal advice model was not trained for.
@@
T2.8 Model supply chain - Data Was publicly available data used? If yes, from where, was data extended or adapted and if yes how? Specify dataset with name and URL, as well as any modifications that the data underwent before being used. For example: Model used MNIST data from ’URL’, where we added 1000 images that were handlabelled by our team.
@@
T2.9 Model Supply Chain - Model weights Are the model weights derived from an existing model, and if yes, which one and how were they adapted? If the original design or weights have a name, provide this name or possibly a URL. Further describe all steps the model undewerwent to adapt it to the current status. For example: Based on Llama3.2:1B, but finetuned on the above described dataset for n iterations using optimzier o and loss l.
@@
T2.10 Model Supply Chain - Code What public code was used to generate and run the model? Provide libraries used with their version name, and also ennumnerate github repositories with link and access data or other re-used code from the project. For example: Pytorch version 2.4.1. and Code from github repo ’this url’ was used, forked January 1st 2024.
@@
T2.11 System prompt If the system is a generative AI system, provide the system prompt. If known, report the entire system, prompt as used in the setting where the incident occurred. For example: ’please be polite to the user and don’t hallucinate. The user writes: ’
@@
Table 3: Reporting scheme with description for annotators and examples, second third (T3.1)-T3.9.
@@
Incident specification T3.1 Materialization status Has the incident been reported prior or after its materialization? Was the attack found during red-teaming or scientific analysis, or as part of an exploited breach? Red teaming
T3.2 Root cause Describe the root cause of the incident. What is the underlying reason, design choice, or property of the system that enables the attack? ’Attacker relied on infinity norm attack the model wasn’t tested for’, or ’API allowed infinite amount of queries, hence enabling crating a copycat training dataset’
T3.3 Malicious Intent Is there evidence that an entity or a person purposely caused the incident by targeting the AI systems security? Is the incident an example of an attempt to destroy, expose, alter, disable, steal or gain unauthorized access to or make unauthorized use of the AI system? ’Yes, the attacker altered the sample until it passed the AI security check’ or ’No, the user did not know that the system was not trained on this specific subset of the training data (Disease B). Hence, the system provided a wrong output.’
T3.4 Attack type Specify the concrete type of attack that has been carried out. Any attack framework, tagging based on the attacks in Sect. 2.4 Poisoning
T3.5 Target asset Did the attack target the model explicitly and the model only? Based on the definition in Sect. 2.2 Yes, model was targeted.
T3.6 Hardware What hardware was used, and could it have affected the incident? Please describe the specific hardware used on the device or server the application ran, and whether the setting implies that the hardware played a role in the incident. Mac M1, could not have affected the incident OR could have affected the incident via a glitch.
T3.7 Access Which access was given to whom for model, inputs, training, test or evaluation data? Specify which levels of access were given to which users regarding each component. Training and evaluation data as well as model was only available to internal model team. The model could be queried via API by the client, but access control prevented any other access by 3rd parties to the model.
T3.8 Security measures What security measures where used? Where they circumvented and if yes how? Give a list of specific security measures that were used including their configurations. Adversarial training using FGSM with epsilon 0.1, 0.2 and 0.3 was used on randomly selected 10% of the dataset were used. Defense was not circumvented as attack was based on L2 norm, not $L_{\infty}$ norm.
T3.9 Fail safe measures Where there any fail-safe measures implemented? Describe any fail-safe measures, such as rebooting the system, not outputting a classification, etc. System fell back on human intervention to prevent further harm.
Table 4: Reporting scheme with description for annotators and examples, last third (T4.1)-T5.2.
@@
Incident Implications T4.1 Regulatory violation List specific regulatory requirements that are violated due to the incident. Please specify any concrete legislation, standard or policy that has been violated by the incidents’ occurrence. European Union AI Act
T4.2 Resulting harm What harm did the incident cause to any participating stakeholder? Did the incident lead to any of the following types of harm? Affecting physical health, inducing financial loss, damaging a physical object, affecting intangible property, affecting infrastructure, affecting natural environment, violation of human rights, civil liberties, civil rights or democratic norms. Harm to physical property of all users, mainly food but also some devices due to malfunction of the fridge AI.
T4.3 Security Impact Describe how the incident affected the confidentiality, integrity, availability of the AI system. Describe whether the incident caused any leakage of confidential data, lead the system to output a wrong result or to exhibit a bias (integrity), or lead to the system being down or otherwise unusable (availability). During the incident, samples were manipulated to be classified as vehicles and thus harmed integrity.
T4.4 Generalization Does the incident potentially affect other AI systems? Incident descriptions should allow to determine for a third model if a vulnerability in that model is likely. Other AI systems might also be affected if relying on the same dataset ”xyz”.
Appendix T5.1 Links to other incidents Links to similar / same cases in other /same database(s) In case you are aware of a incidents that has been reported in another database or that is similar to this incident in this database This incident has had similar consequences like AIID incidents 38 or simply ’URL’
T5.2 Comments Additional comments. Anything you’d like to add which has not been covered in the previous fields. The attack does not work on model ABC1.2 and ABC1.4, but only on ABC1.3